# Upho lattices II: ways of realizing a core

### v1

@@ 3. Colorable upho posets, automorphisms, and finiteness
First, suppose that $y^{\prime}$ and $z^{\prime}$ are covered in common by some element $w\coloneq y^{\prime}\vee z^{\prime}$, i.e., that $\rho(y^{\prime}\vee z^{\prime})=k$. Since $\rho(w)=k$, $\mathcal{L}$ agrees with $\mathcal{F}_{n}$ through rank $k$, and $w$ covers more than one element, $w$ is a flip of some vertex two ranks below it, and all the elements covered by $w$ share a single lower cover. Therefore $y^{\prime}\wedge z^{\prime}$ has rank $k-2>0$. However, in this case the elements $x,y,z$ all belong to $V_{y^{\prime}\wedge z^{\prime}}\cong\mathcal{L}$, and have rank $\leq 3$ in this poset; this contradicts the hypothesis that $\mathcal{L}$ and $\mathcal{F}_{n}$ agree to rank $k$. Therefore this case does not occur.
Alternatively, it must be the case that for every possible choice of $x,y,z,y^{\prime},z^{\prime}$ in $\mathcal{L}$ such that $y^{\prime}\lessdot y\lessdot x\gtrdot z\gtrdot z^{\prime}$ for which $y$ and $z$ do not cover anything in common, there is no element that covers both $y^{\prime}$ and $z^{\prime}$. For $i\geq 0$, let $a_{i}$ be the size of the $i$th rank of $\mathcal{F}_{n}$. By Corollary 2.3, this is also the size of the $i$th rank of $\mathcal{L}$. Moreover, these numbers satisfy the recurrence relation $a_{i}=na_{i-1}-(n-1)a_{i-2}$ for all $i\geq 2$. Now let us count cover relations between ranks $k$ and $k+1$, which we refer to as “edges,” because they are edges in the Hasse diagram of $\mathcal{L}$. On one hand, since $\mathcal{L}$ is upho with core $M_{n}$, there are $n\cdot a_{k}$ of these edges. On the other hand, it follows from the defining hypothesis of this case that $n\cdot a_{k-1}$ of these edges belong to copies of $M_{n}$ with minimum elements at rank $k-1$ (i.e., all those subgraphs of the Hasse diagram are edge-disjoint), so that there are $n\cdot a_{k}-n\cdot a_{k-1}=a_{k+1}-a_{k-1}$ other edges. Observe also that there are at most $a_{k-1}$ elements at rank $k+1$ that are the maximum elements of the copies of $M_{n}$ whose minimum elements lie at rank $k-1$, and strictly fewer than $a_{k-1}$ if some choice of $y$ and $z$ lie in two of these copies of $M_{n}$. Therefore, there are at least $a_{k+1}-a_{k-1}$ “other” elements, and again strictly more under the same condition. If the inequalities are strict, this leads to a contradiction because there are more than $a_{k+1}-a_{k-1}$ “other” elements at rank $k+1$, each of which covers at least one element, but only $a_{k+1}-a_{k-1}$ edges to connect them to. If the inequality is actually equality, then we instead get a contradiction because the $x$ corresponding to our choice of $y$ and $z$, which covers at least two elements, must be among the “other” elements, and so there are still not enough edges available.
@@
Choosing any $f\colon[n]\to[n]$ which has image of size $1$ yields the dominating vertex construction $\mathcal{D}_{n}$ for $\mathcal{L}(f)$. To see this, for convenience consider the function $f$ defined by $f(1)=f(2)=\cdots=f(n)=n$. Let $S=\{s_{1},\ldots,s_{n}\}$ and again let $S^{*}$ denote the set of words over the alphabet $S$. Then, for any two words $v,w\in S^{*}$ of the same length, we have $vs_{n}=ws_{n}$ in $M(f)$, while $vs_{i}$ and $ws_{j}$ will be distinct, and be distinct from $vs_{n}$, for all $11\leq i,j\leq n-$ 1. Thus, $\mathcal{L}(f)$ is indeed $\mathcal{D}_{n}$.
On the other hand, choosing any bijective $f\colon[n]\to[n]$ yields the flip construction $\mathcal{F}_{n}$ for $\mathcal{L}(f)$. To see this, for convenience consider the identity function $f$ on $[n]$. By the symmetry with respect to left and right multiplication of the defining relations for this $f$, we have that $M(f)$ is not just left-cancellative, but also right-cancellative. So now suppose that $v,w\in M(f)$ are two elements which (when viewed as words in $S^{*}$) have length $k$, and which satisfy $vs_{i}=ws_{j}$ for some $1\leq i,j\leq n$. There are two possibilities. If $i=j$, then by right-cancellativity we have $v=w$. If $i\neq j$, then we claim that there is $u\in M(f)$ of length $k-1$ such that $v=us_{i}$ and $w=us_{j}$. Indeed, to convert the final $s_{i}$ in $vs_{i}$ to an $s_{j}$, we must have $vs_{i}=us_{i}s_{i}=us_{j}s_{j}$ for some $u\in M(f)$ of length $k-1$. Then since $ws_{j}=vs_{i}=us_{j}s_{j}$, by right-cancellativity again we get $v=us_{i}$ and $w=us_{j}$, as claimed. Thus two elements of the same rank in $\mathcal{L}(f)$ have a common cover exactly when they cover something in common, and so $\mathcal{L}(f)$ is indeed $\mathcal{F}_{n}$. ∎
@@
To prove that the monoid $M$ is left-cancellative, a standard inductive argument implies it is enough to show that whenever $sb=sc$ for $b,c\in M$ arbitrary elements and $s\in S$ a generator, then we have $b=c$. So suppose that $b,c\in M$ and $s\in S$ satisfy $sb=sc$. Because the relations defining $M$ is homogeneous, it must be that $b$ and $c$ (when viewed as words in $S^{*}$) have the same length $n\geq 0$. We will prove that $b=c$ by induction on $n\geq 0$. If $n=0$ the claim is trivial, so suppose that $n\geq 1$ and that the claim has been proved for smaller values of $n$. By the same standard inductive argument, knowing the claim for smaller values of $n$ also means that whenever $ab=ac$ for any elements $a,b,c\in M$ for which the length of $ab$ (which must be the length of $ac$) is less than $n$, we have $b=c$.
By supposition, we can convert $sb$ (viewed as a word in $S^{*}$) to $sc$ by applying a series of the relations in $R$. If when applying these relations, we never change the first letter, then clearly $b=c$ in $M$. So suppose that we do change the first letter at some point. For a generator $t\in S$, let $w_{t}$ denote the unique word (if it exists) for which $tw_{t}$ appears in the relations in $R$. Then set $s_{0}\coloneqq s$, and suppose that as we convert $sb$ to $sc$, the first time we change the first letter is $s_{0}w_{s_{0}}v_{0}=s_{1}w_{s_{1}}v_{0}$ for some other generator $s_{1}\in S$ and some word $v_{0}\in S^{*}$. Notice that we have $b=w_{s_{0}}v_{0}$ in $M$ because by supposition we never modified the first letter before getting to $s_{0}w_{s_{0}}v_{0}$. Then suppose the next time we change the first letter, it is $s_{1}w_{s_{1}}v_{1}=s_{2}w_{s_{2}}v_{1}$ for some $s_{2}\in S$, $v_{1}\in S^{*}$. The same reasoning as before says that $w_{s_{1}}v_{0}=w_{s_{1}}v_{1}$ in $M$. By the inductive hypothesis, because $w_{s_{1}}v_{0}$ has length $n-1$, this actually means that $v_{0}=v_{1}$ in $M$. We continue defining $s_{i}\in S$ and $v_{i}\in S^{*}$ in the same fashion, until the last time we change the first letter, say to $s_{k}w_{s_{k}}v_{k-1}$. By the same inductive argument repeated $k-1$ times, we have $v_{0}=v_{1}=\cdots=v_{k-1}$ in $M$. Notice also that we must have $s_{k}=s$, and that $w_{s_{k}}v_{k-1}=c$ in $M$ because after $s_{k}w_{s_{k}}v_{k-1}$ we never change the first letter. But this means that $b=w_{s}v_{0}=w_{s}v_{k-1}=c$, as claimed. ∎
@@
Let $\mathcal{P}$ be a finite-type $\mathbb{N}$-graded poset. Suppose that for any $x,y\in\mathcal{P}$, $x$ and $y$ have an upper bound, and moreover, if $x$ and $y$ both cover an element $z$, then they have a least upper bound $x\vee y$. Then $\mathcal{P}$ is a lattice.
@@
As indicated, this is essentially a special case of the famous “BEZ lemma” [2, Lemma 2.1]. Let $\mathcal{P}$ be as in the statement of the lemma. We first show that for any $x,y\in\mathcal{P}$, their join $x\vee y$ exists in $\mathcal{P}$. By assumption, an upper bound for $x$ and $y$ exists, call it $z\in\mathcal{P}$. Now, $[\hat{0},z]$ is a finite poset, and it has the property that for any pair of elements covering a common element, their join exists, so the BEZ lemma applies and says that $[\hat{0},z]$ is a lattice. In particular, a join of $x$ and $y$ exists in $[\hat{0},z]$, call it $w$. Now, we claim this $w$ is actually the least upper bound of $x$ and $y$ in $\mathcal{P}$ as well. Indeed, because $[\hat{0},z]$ is an interval, it is clear that there cannot be an upper bound in $\mathcal{P}$ for $x$ and $y$ that is smaller than $w$. And so if $z^{\prime}\in\mathcal{P}$ is any upper bound for $x$ and $y$, then, letting $z^{\prime\prime}$ be an upper bound for $w$ and $z^{\prime}$, the BEZ lemma applies also to $[\hat{0},z^{\prime\prime}]$ to show that indeed $w\leq z^{\prime}$.
Having shown that joins of all pairs of elements exist, it is easy to show that meets exist, i.e., that $\mathcal{P}$ is a lattice. Let $x,y\in\mathcal{P}$. Then $[\hat{0},x\vee y]$ is a finite join semilattice with a minimum, hence is a lattice by [13, Proposition 3.3.1]. So the meet $x\wedge y$ exists in $[\hat{0},x\vee y]$, and this must be their meet in $\mathcal{P}$ as well. ∎
@@
The suffix of this word of length $22k-$ 2 is of the form described by the claim, so by induction it is equal to
@@
It is well known that the partition function $p(n)$ grows super-polynomially. Computational evidence suggests that $\kappa(M_{n})$ should in fact grow at least exponentially. In particular, by considering just the finite portion of the posets up to rank $4$, we found by a brute-force computation in Sage [12] that for $n=3,4,5,6$, the numbers of isomorphism classes of posets of the form $\mathcal{L}(f)$ for functions $f\colon[n]\to[n]$ are at least $4,8,16$, and $35$.
@@ 5. Listing all the ways a finite lattice can be a core
First of all, we note that for $L$ to be a core of an upho lattice, an obvious requirement is that it be a finite graded lattice for which its maximum $\hat{1}$ is the join of its atoms. Hence, we will only consider $L$ of this form.
Let $L$ be a finite graded lattice for which $\hat{1}$ is the join of the atoms. A pre-upho coloring of $L$ is a function $c$ that maps each cover relation of $L$ to an atom of $L$, such that:
@@
  • for each $x\in L\setminus\{\hat{0},\hat{1}\}$, letting $y_{1},\ldots,y_{k}$ be the covers of $x$, there is a rank- and color-preserving embedding of the interval $[x,y_{1}\vee\cdots\vee y_{k}]$ into $L$.
(By an embedding of $P$ into $Q$ we mean a map $\varphi\colon P\to Q$ which is an isomorphism onto its image. That this isomorphism is rank-preserving means $\rho(\varphi(p))=\rho(p)$ for all $p\in P$; that it is color-preserving means $c(\varphi(x)\lessdot\varphi(y))=c(x\lessdot y)$ for all $x\lessdot y\in P$.) Note that a rank-preserving embedding of $[x,y_{1}\vee\cdots\vee y_{k}]$ into $L$ must in particular send $x$ to $\hat{0}$ and $y_{1},\ldots,y_{k}$ to atoms of $L$.
@@
Let $x\in L\setminus\{\hat{0},\hat{1}\}$ and let $y_{1},\ldots,y_{k}\in L$ be the elements covering $x$. Let $\varphi_{x}\colon V_{x}\to\mathcal{L}$ be an isomorphism verifying that the coloring $c$ of $\mathcal{L}$ is upho. Then $\varphi_{x}$ restricts to a rank- and color-preserving embedding of $[x,y_{1}\vee\cdots\vee y_{k}]\subseteq L$ into $L$. Indeed, the only thing that needs to be checked is that $\varphi_{x}([x,y_{1}\vee\cdots\vee y_{k}])\subseteq L$, because that $\varphi_{x}$ restricts to a rank- and color-preserving embedding of $[x,y_{1}\vee\cdots\vee y_{k}]$ into $\mathcal{L}$ is clear. To see that $\varphi_{x}([x,y_{1}\vee\cdots\vee y_{k}])\subseteq L$, first note that since the core $L$ is an interval in $\mathcal{L}$, it is a sublattice of $\mathcal{L}$. Thus, the join $y_{1}\vee\cdots\vee y_{k}$ is the same whether considered in $L$ or $\mathcal{L}$. Since $y_{1}\vee\cdots\vee y_{k}\in\mathcal{L}$ is less than the join of all the elements in $\mathcal{L}$ covering $x$ (whose image under $\varphi_{x}$ is the maximum of the core $L$), it must be that $\varphi_{x}(y_{1}\vee\cdots\vee y_{k})$, and hence all of $\varphi_{x}([x,y_{1}\vee\cdots\vee y_{k}])$, belongs to $L$. ∎
@@
[…]
@@
First of all, the conjecture only guarantees that there is a rank-preserving embedding of $L$ into $\mathcal{L}$. This means that $L$ sits inside the core of $\mathcal{L}$, but the core could potentially be bigger than just $L$. Indeed, Figure 5 shows an example where the core of the output $\mathcal{L}$ is strictly bigger than the input $L$. But this is not such a serious problem for our desired algorithm, as we can check that the core of any output $\mathcal{L}$ is really $L$ just by looking at a finite portion of $\mathcal{L}$.

### v2

@@ 3. Colorable upho posets, automorphisms, and finiteness
First, suppose that $y^{\prime}$ and $z^{\prime}$ are covered in common by some element $w\coloneqq y^{\prime}\vee z^{\prime}$, i.e., that $\rho(y^{\prime}\vee z^{\prime})=k$. Since $\rho(w)=k$, $\mathcal{L}$ agrees with $\mathcal{F}_{n}$ through rank $k$, and $w$ covers more than one element, $w$ is a flip of some vertex two ranks below it, and all the elements covered by $w$ share a single lower cover. Therefore $y^{\prime}\wedge z^{\prime}$ has rank k2>0k-2>0. However, in this case the elements $x,y,z$ all belong to $V_{y^{\prime}\wedge z^{\prime}}\cong\mathcal{L}$, and have rank $\leq 3$ in this poset; this contradicts the hypothesis that $\mathcal{L}$ and $\mathcal{F}_{n}$ agree to rank $k$. Therefore this case does not occur.
Alternatively, it must be the case that for every possible choice of $x,y,z,y^{\prime},z^{\prime}$ in $\mathcal{L}$ such that $y^{\prime}\lessdot y\lessdot x\gtrdot z\gtrdot z^{\prime}$ for which $y$ and $z$ do not cover anything in common, there is no element that covers both $y^{\prime}$ and $z^{\prime}$. For $i\geq 0$, let $a_{i}$ be the size of the $i$th rank of $\mathcal{F}_{n}$. By Corollary 2.3, this is also the size of the $i$th rank of $\mathcal{L}$. Moreover, these numbers satisfy the recurrence relation $a_{i}=na_{i-1}-(n-1)a_{i-2}$ for all $i\geq 2$. Now let us count cover relations between ranks $k$ and $k+1$, which we refer to as “edges,” because they are edges in the Hasse diagram of $\mathcal{L}$. On one hand, since $\mathcal{L}$ is upho with core $M_{n}$, there are $n\cdot a_{k}$ of these edges. On the other hand, it follows from the defining hypothesis of this case that $n\cdot a_{k-1}$ of these edges belong to copies of $M_{n}$ with minimum elements at rank k1k-1 (i.e., all those subgraphs of the Hasse diagram are edge-disjoint), so that there are $n\cdot a_{k}-n\cdot a_{k-1}=a_{k+1}-a_{k-1}$ other edges. Observe also that there are at most $a_{k-1}$ elements at rank $k+1$ that are the maximum elements of the copies of $M_{n}$ whose minimum elements lie at rank k1k-1, and strictly fewer than $a_{k-1}$ if some choice of $y$ and $z$ lie in two of these copies of $M_{n}$. Therefore, there are at least $a_{k+1}-a_{k-1}$ “other” elements, and again strictly more under the same condition. If the inequalities are strict, this leads to a contradiction because there are more than $a_{k+1}-a_{k-1}$ “other” elements at rank $k+1$, each of which covers at least one element, but only $a_{k+1}-a_{k-1}$ edges to connect them to. If the inequality is actually equality, then we instead get a contradiction because the $x$ corresponding to our choice of $y$ and $z$, which covers at least two elements, must be among the “other” elements, and so there are still not enough edges available.
@@
Choosing any $f\colon[n]\to[n]$ which has image of size $1$ yields the dominating vertex construction $\mathcal{D}_{n}$ for $\mathcal{L}(f)$. To see this, for convenience consider the function $f$ defined by $f(1)=f(2)=\cdots=f(n)=n$. Let $S=\{s_{1},\ldots,s_{n}\}$ and again let $S^{*}$ denote the set of words over the alphabet $S$. Then, for any two words $v,w\in S^{*}$ of the same length, we have $vs_{n}=ws_{n}$ in $M(f)$, while $vs_{i}$ and $ws_{j}$ will be distinct, and be distinct from $vs_{n}$, for all $1\leq i,j\leq n-1$. Thus, $\mathcal{L}(f)$ is indeed $\mathcal{D}_{n}$.
On the other hand, choosing any bijective $f\colon[n]\to[n]$ yields the flip construction $\mathcal{F}_{n}$ for $\mathcal{L}(f)$. To see this, for convenience consider the identity function $f$ on $[n]$. By the symmetry with respect to left and right multiplication of the defining relations for this $f$, we have that $M(f)$ is not just left-cancellative, but also right-cancellative. So now suppose that $v,w\in M(f)$ are two elements which (when viewed as words in $S^{*}$) have length $k$, and which satisfy $vs_{i}=ws_{j}$ for some $1\leq i,j\leq n$. There are two possibilities. If $i=j$, then by right-cancellativity we have $v=w$. If $i\neq j$, then we claim that there is $u\in M(f)$ of length k1k-1 such that $v=us_{i}$ and $w=us_{j}$. Indeed, to convert the final $s_{i}$ in $vs_{i}$ to an $s_{j}$, we must have $vs_{i}=us_{i}s_{i}=us_{j}s_{j}$ for some $u\in M(f)$ of length k1k-1. Then since $ws_{j}=vs_{i}=us_{j}s_{j}$, by right-cancellativity again we get $v=us_{i}$ and $w=us_{j}$, as claimed. Thus two elements of the same rank in $\mathcal{L}(f)$ have a common cover exactly when they cover something in common, and so $\mathcal{L}(f)$ is indeed $\mathcal{F}_{n}$. ∎
@@
To prove that the monoid $M$ is left-cancellative, a standard inductive argument implies it is enough to show that whenever $sb=sc$ for $b,c\in M$ arbitrary elements and $s\in S$ a generator, then we have $b=c$. So suppose that $b,c\in M$ and $s\in S$ satisfy $sb=sc$. Because the relations defining $M$ are homogeneous, it must be that $b$ and $c$ (when viewed as words in $S^{*}$) have the same length $n\geq 0$. We will prove that $b=c$ by induction on $n\geq 0$. If $n=0$ the claim is trivial, so suppose that $n\geq 1$ and that the claim has been proved for smaller values of $n$. By the same standard inductive argument, knowing the claim for smaller values of $n$ also means that whenever $ab=ac$ for any elements $a,b,c\in M$ for which the length of $ab$ (which must be the length of $ac$) is less than $n$, we have $b=c$.
By supposition, we can convert $sb$ (viewed as a word in $S^{*}$) to $sc$ by applying a series of the relations in $R$. If when applying these relations, we never change the first letter, then clearly $b=c$ in $M$. So suppose that we do change the first letter at some point. For a generator $t\in S$, let $w_{t}$ denote the unique word (if it exists) for which $tw_{t}$ appears in the relations in $R$. Then set $s_{0}\coloneqq s$, and suppose that as we convert $sb$ to $sc$, the first time we change the first letter is $s_{0}w_{s_{0}}v_{0}=s_{1}w_{s_{1}}v_{0}$ for some other generator $s_{1}\in S$ and some word $v_{0}\in S^{*}$. Notice that we have $b=w_{s_{0}}v_{0}$ in $M$ because by supposition we never modified the first letter before getting to $s_{0}w_{s_{0}}v_{0}$. Then suppose the next time we change the first letter, it is $s_{1}w_{s_{1}}v_{1}=s_{2}w_{s_{2}}v_{1}$ for some $s_{2}\in S$, $v_{1}\in S^{*}$. The same reasoning as before says that $w_{s_{1}}v_{0}=w_{s_{1}}v_{1}$ in $M$. By the inductive hypothesis, because $w_{s_{1}}v_{0}$ has length n1n-1, this actually means that $v_{0}=v_{1}$ in $M$. We continue defining $s_{i}\in S$ and $v_{i}\in S^{*}$ in the same fashion, until the last time we change the first letter, say to $s_{k}w_{s_{k}}v_{k-1}$. By the same inductive argument repeated k1k-1 times, we have $v_{0}=v_{1}=\cdots=v_{k-1}$ in $M$. Notice also that we must have $s_{k}=s$, and that $w_{s_{k}}v_{k-1}=c$ in $M$ because after $s_{k}w_{s_{k}}v_{k-1}$ we never change the first letter. But this means that $b=w_{s}v_{0}=w_{s}v_{k-1}=c$, as claimed. ∎
@@
Let $\mathcal{P}$ be a finite-type $\mathbb{N}$-graded poset. Suppose that for any $y\in\mathcal{P}, xx$ and $y$ have an upper bound, and moreover, if $x$ and $y$ both cover an element $z$, then they have a least upper bound $x\vee y$. Then $\mathcal{P}$ is a lattice.
@@
As indicated, this is essentially a special case of the famous “BEZ lemma” [2, Lemma 2.1]. Let $\mathcal{P}$ be as in the statement of the lemma. We first show that for any $x,y\in\mathcal{P}$, their join $x\vee y$ exists in $\mathcal{P}$. By assumption, an upper bound for $x$ and $y$ exists, call it $z\in\mathcal{P}$. Now, $\hat{0},z]$ is a finite poset, and it has the property that for any pair of elements covering a common element, their join exists, so the BEZ lemma applies and says that $\hat{0},z]$ is a lattice. In particular, a join of $x$ and $y$ exists in $\hat{0},z]$, call it $w$. Now, we claim this $w$ is actually the least upper bound of $x$ and $y$ in $\mathcal{P}$ as well. Indeed, because $\hat{0},z]$ is an interval, it is clear that there cannot be an upper bound in $\mathcal{P}$ for $x$ and $y$ that is smaller than $w$. And so if $z^{\prime}\in\mathcal{P}$ is any upper bound for $x$ and $y$, then, letting $z^{\prime\prime}$ be an upper bound for $w$ and $z^{\prime}$, the BEZ lemma applies also to $\hat{0},z^{\prime\prime}]$ to show that indeed $w\leq z^{\prime}$.
Having shown that joins of all pairs of elements exist, it is easy to show that meets exist, i.e., that $\mathcal{P}$ is a lattice. Let $x,y\in\mathcal{P}$. Then $\hat{0},x\vee y]$ is a finite join semilattice with a minimum, hence is a lattice by [13, Proposition 3.3.1]. So the meet $x\wedge y$ exists in $\hat{0},x\vee y]$, and this must be their meet in $\mathcal{P}$ as well. ∎
@@
The suffix of this word of length 2k22k-2 is of the form described by the claim, so by induction it is equal to
@@
After seeing the first version of this paper, David Speyer explained to us how our methods can be extended to show that $\kappa(M_{n})\geq F_{n-3}$ for all $n\geq 3$, where $F_{n}$ is the $n$th Fibonacci number. (Recall that the Fibonacci numbers are defined recursively by $F_{0}=0$, $F_{1}=1$, and $F_{n}=F_{n-1}+F_{n-2}$ for $n\geq 2$. Unlike the partition numbers, which grow super-polynomially but sub-exponentially, the Fibonacci numbers grow exponentially.) We thank him for allowing us to include a sketch of his argument here. Given a composition $\alpha=(\alpha_{1},\ldots,\alpha_{k})$ of n2n-2 into parts of size at least $2$, define the function $f_{\alpha}\colon[n]\to[n]$ as follows:
  • $f_{\alpha}(1)=1$;
  • $f_{\alpha}(i)=i-1$ for $2\leq i\leq 2k+2$;
  • if $2(k-j+2)+(\alpha_{1}+\alpha_{2}+\cdots+\alpha_{j-1})<i\leq 2(k-j+1)+(\alpha_{1}+\alpha_{2}+\cdots+\alpha_{j})$ for $1\leq j\leq k$, then $f_{\alpha}(i)=2j+1$.
Then one can show that for two such compositions $\alpha$ and $\beta$ of n2n-2, the upho lattices $\mathcal{L}(f_{\alpha})$ and $\mathcal{L}(f_{\beta})$ are isomorphic if and only if $\alpha=\beta$. And for $n\geq 3$, there are $F_{n-3}$ such compositions $\alpha$ of n2n-2.
@@ 5. Listing all the ways a finite lattice can be a core
First of all, we note that for $L$ to be a core of an upho lattice, an obvious requirement is that it be a finite graded lattice whose maximum element $\hat{1}$ is the join of its atoms. Hence, we will only consider $L$ of this form.
Let $L$ be a finite graded lattice for which $\hat{1}$ is the join of the atoms. A pre-upho coloring [^5] Our notion of pre-uhpo coloring of a finite graded lattice is similar to the notion of “colored semi-upho poset” considered in [6]. of $L$ is a function $c$ that maps each cover relation of $L$ to an atom of $L$, such that:
@@
  • for each $x\in L\setminus\{\hat{0},\hat{1}\}$, letting $y_{1},\ldots,y_{k}$ be the covers of $x$, there is a rank- and color-preserving embedding of the interval $x,y_{1}\vee\cdots\vee y_{k}]$ into $L$.
(By an embedding of $P$ into $Q$ we mean a map $\varphi\colon P\to Q$ which is an isomorphism onto its image. That this isomorphism is rank-preserving means $\rho(\varphi(p))=\rho(p)$ for all $p\in P$; that it is color-preserving means $c(\varphi(x)\lessdot\varphi(y))=c(x\lessdot y)$ for all $x\lessdot y\in P$.) Note that a rank-preserving embedding of $x,y_{1}\vee\cdots\vee y_{k}]$ into $L$ must in particular send $x$ to $\hat{0}$ and $y_{1},\ldots,y_{k}$ to atoms of $L$.
@@
Let $x\in L\setminus\{\hat{0},\hat{1}\}$ and let $y_{1},\ldots,y_{k}\in L$ be the elements covering $x$. Let $\varphi_{x}\colon V_{x}\to\mathcal{L}$ be an isomorphism verifying that the coloring $c$ of $\mathcal{L}$ is upho. Then $\varphi_{x}$ restricts to a rank- and color-preserving embedding of $x,y_{1}\vee\cdots\vee y_{k}]\subseteq L$ into $L$. Indeed, the only thing that needs to be checked is that $\varphi_{x}([x,y_{1}\vee\cdots\vee y_{k}])\subseteq L$, because that $\varphi_{x}$ restricts to a rank- and color-preserving embedding of $x,y_{1}\vee\cdots\vee y_{k}]$ into $\mathcal{L}$ is clear. To see that $\varphi_{x}([x,y_{1}\vee\cdots\vee y_{k}])\subseteq L$, first note that since the core $L$ is an interval in $\mathcal{L}$, it is a sublattice of $\mathcal{L}$. Thus, the join $y_{1}\vee\cdots\vee y_{k}$ is the same whether considered in $L$ or $\mathcal{L}$. Since $y_{1}\vee\cdots\vee y_{k}\in\mathcal{L}$ is less than the join of all the elements in $\mathcal{L}$ covering $x$ (whose image under $\varphi_{x}$ is the maximum of the core $L$), it must be that $\varphi_{x}(y_{1}\vee\cdots\vee y_{k})$, and hence all of $\varphi_{x}([x,y_{1}\vee\cdots\vee y_{k}])$, belongs to $L$. ∎
@@
$M=\left\langle s_{1},\ldots,s_{r}\mid\parbox{271.0125pt}{\begin{center} $c(\hat{0}=x_{0}\lessdot x_{1})c(x_{1}\lessdot x_{2})\cdots c(x_{k-1}\lessdot x_{k}=x_{1}\vee y_{1})=$ \\ $c(\hat{0}=y_{0}\lessdot y_{1})c(y_{1}\lessdot y_{2})\cdots c(y_{k-1}\lessdot y_{k}=x_{1}\vee y_{1})$ \end{center}}\right\rangle$
@@
First of all, the conjecture only guarantees that there is a rank-preserving embedding of $L$ into $\mathcal{L}$. This means that $L$ sits inside the core of $\mathcal{L}$, but the core could potentially be bigger than just $L$. Indeed, Figure 5 shows an example where the core of the output $\mathcal{L}$ is strictly bigger than the input $L$. But this is not such a serious problem for our desired algorithm, as we can check that the core of any output $\mathcal{L}$ is really exactly $L$ just by looking at a finite portion of $\mathcal{L}$.